OpenStudy (anonymous):

Can someone please explain the steps to get from -((9a)^1/2) to 3^-1a^-1/2

OpenStudy (anonymous):

\[ -(9a)^{1\over2}=-\sqrt{9a}=-\sqrt{9}\times\sqrt{a}=-\sqrt{3^2}\times a^{1\over2} =-3 a^{1\over2} \]

OpenStudy (anonymous):

did you write the question correctly?

OpenStudy (anonymous):

Ok, and that can be taken one step further to \[3^{-1}a ^{-_{2}^{1}}\]

OpenStudy (anonymous):

I'm pretty sure I did. I'm simplifying an exercise in my text book, and that is the answer, I was just unsure how they got to it. + and - mixed with exponents and radicals still confuse me

OpenStudy (anonymous):

nope.. remember\[a^{-m}={1\over a^m}\] so, the given way, it cannot happen. (unless more information about "a" or something is provided)

OpenStudy (anonymous):

by the above rule, we can write it as: \[-3a^{1\over2}=\frac{-1}{3^{-1}a^{-1\over2}}\]

OpenStudy (anonymous):

ok, what about if I forgot to tell you that this part of the question was the denominator and the initial - is because it was moved up to the numerator?

OpenStudy (anonymous):

that sort of fits with the rule you just showed me, doesn't it?

OpenStudy (anonymous):

SEE!!! math dont lie

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

sheeesh, I don't think I'll ever get this sorted in my head...

OpenStudy (anonymous):

Thank you for your help :D

OpenStudy (anonymous):

yer welcome <:)